# Generative Diffusion Model Driven Massive Random Access in Massive MIMO Systems

### v2

@@ Appendix A Structure of the Heterogeneous Transformer Network
Fig. 17 illustrates the structure of the intermediate layer in the proposed VPL-AUDNet network. First, for the input pilot feature sequences $\{\bm{z}_{k}^{0}\}_{k=0}^{K}$, and the received signal feature sequence $\bm{z}_{K+1}^{0}$, two distinct linear layers are employed for linear embedding. This process results in the sequences, $\{\bm{z}_{k}^{1}\}_{k=1}^{K}$, with dimensions $(B,K+1,d)$, where $B$ is the batch size, $K+1$ is the sequence length, and $d$ is the feature dimension. Subsequently, the sequence undergoes processing through $N$ alternating heterogeneous Transformer encoder layers and PLAM networks, and finally produce $\{\bm{z}_{k}^{N+1}\}_{k=1}^{K}$ for the output layer of the VPL-AUDNet. Fig. 17 also demonstrates that the unique aspect of the heterogeneous Transformer network, in contrast to the traditional Transformer, is its use of distinct linear matrix weights for input embedding, MHA, and FFN components.

### v3

@@ Appendix A Structure of the Heterogeneous Transformer Network
Fig. 17 illustrates the structure of the intermediate layer in the proposed VPL-AUDNet network. First, for the input pilot feature sequences $\{\bm{z}_{k}^{0}\}_{k=0}^{K}$, and the received signal feature sequence $\bm{z}_{K+1}^{0}$, two distinct linear layers are employed for linear embedding. This process results in the sequences, $\{\bm{z}_{k}^{1}\}_{k=1}^{K}$, with dimensions $(B,K+1,d)$, where $B$ is the batch size, $K+1$ is the sequence length, and $d$ is the feature dimension. Subsequently, the sequence undergoes processing through NN alternating heterogeneous Transformer encoder layers and PLAM networks, and finally produce $\{\bm{z}_{k}^{N+1}\}_{k=1}^{K}$ for the output layer of the VPL-AUDNet. Fig. 17 also demonstrates that the unique aspect of the heterogeneous Transformer network, in contrast to the traditional Transformer, is its use of distinct linear matrix weights for input embedding, MHA, and FFN components.